Show complete solution on the upper half of the graphing paper and the graph on the lower half portion. One function is the one graphing paper.

1.f(x)=2 sin (x-​π/​4)

2.f(x)=sin (x+π/​3)

3.f(x)=3 cos (x+π/​6)

4.f(x)=2 cos (x-2π)

5.f(x)=tan x


1
Expert's answer
2021-01-27T10:03:57-0500

Q1.

f(x) = 2sin(x - π/4)

=> f(x)= 2(sinx cosπ/4 - cosx sinπ/4)

=> f(x) = 2((1/"\sqrt{2}" ) sinx - (1/"\sqrt{2}" )cosx)

=> f(x) = "\sqrt{2}" (sinx - cosx)

Graph of y = sinx will be shifted right by π/4 and then it will be stretched twice vertically



Q2.

f(x) = sin(x + π/3)

=> f(x)= (sinx cosπ/3 + cosx sinπ/3)

=> f(x) = ((1/2) sinx + ("\sqrt{3}" /2)cosx)

=> f(x) = (1/2)(sinx +"\sqrt{3}" cosx)

Graph of y = sinx will be shifted left by π/3



Q3.

f(x) = 3cos(x + π/6)

f(x) = 3{ cosx cosπ/6 - sinx sin π/6}

=> f(x) = 3{("\sqrt{3}" /2)cosx - (1/2)sinx}

=> f(x) = (3/2)("\sqrt{3}" cosx - sinx)

Graph of y = cosx will be shifted left by π/6 and then it will be stretched three times vertically



Q4.

f(x) = 2cos(x - 2π)

=> f(x) = 2cos(2π - x)

=> f(x) = 2cosx

Graph of y = cosx will be stretched twice vertically.




Q5.

f(x) = tanx

f(x) = tanx is a discontinuous function at x = (2n+1)(π/2), n is integer. tanx is increasing function from -∞ to ∞ as x changes from -π/2 to π/2 . Similarly in the other slots like (-5π/2, -3π/2),(-3π/2, -π/2) (π/2, 3π/2) , (3π/2, 5π/2) etc.